AO Commissioning on 11 June 2011

Observer(s): D. Miller, JC. Guerra, K. Boutsia, E. Pinna, L Busoni, A. Puglisi, S. Esposito, J. Hill, R. Ragazzoni (LBTO)

Telescope Operator: S. Allanson (LBTO)

SW Support: N. Cushing (Tucson)

Telescope Support: J. Urban (LBTO)

Instrument Support: T. Shih(LBTO)

Summary

DIMM seeing reported to be form 0.9" to 1.2" all night.

Filter Wheel

In the afternoon, a discussion about the filter wheel resulted in two software modification to the filter wheel software by Craig.

  • First, Craig slowed the stepper motor down (by a factor of two I believe) in case the problem in positioning the filter wheel lost step because of trying to move too fast. Craig did say he was turning the motor at the manufacturer suggested speed, which is quite a bit below the stall(?) speed for this motor (again, from the manufacturer). However, it was worth examining if a slower motor would position the filter wheel more reliably.
  • Second, Craig modified the filter wheel software to allow floating point positioning of the filter wheel, verses the previous integer only. What this means is that the user can ask to filter to move to position 5.2 rather than to just filter 5 at position 5. In this way, if the filter is found to not be positioned with a filter centered by reading off the motor encoder values on top of the motor, the filter could be nudged a bit (a fraction of a filter position) to then get it centered. Previously if we found the filter wheel out of position, the only recourse was to re-home the filter, as for the integer filter position again and examine the motor encoder count to see if the filter was properly positioned.

A webcam was mounted so the filter wheel motor encoder could be viewed while operating the filter wheel to confirm its position. The filter wheel failed to properly position itself only once (Arcetri guy need to confirm this) during the night, which include 30 or so filter changes. It is not clear if the slowing of the stepper motor was the cause for this improvement, or there were other conditions (?) that improved that allowed the filter wheel to position itself properly.

PISCES Characterization

JC and KB took some flat frames in H and Ks at Dust

About 40 minutes was spent during the night observing photometric standards in several bands (J, Ks, H, H2)

Open Loop Offsets

We had tested open loop offsets the previous night, with success, and tonight we used open loop offsets extensively (~40 times) with only a couple of failed offsets. John H. correctly pointed out that in Closed AO Loop the collimation lookup table is not being applied, so when an offset is performed, by default and design, the collimation lookup table is applied. This may cause a slight shift of the field, including the AO Reference star, such that the reference star is out of the Pyramid's capture range after the offset, and the loop does not resume. John made IT #3435. The short term solution may be for Doug to write an IDL script that sends a StepFocus 0.0 every 10 to 15 seconds to cause a collimation lookup table application.

Off-axis AO Reference Star

At the end of the night, when the sky was getting too bright to work on dim star AO characterization, a presetTelescope was sent in ACE mode using an off-axis AO Reference star, 10" away from the Target (on-axis) star. The positioning of the W-Unit stage was good and acquisition and closed AO loop proceeded successfully. Image quality data were collected with the Target and Reference stars both visible on the PISCES chip. The image quality of the Reference star was quite good, and a bit poorer IQ was seen with the Target star (as expected), the the image still looked quite good.

We will still need to send several more presets with off-axis AO Reference stars, along with some offsets when off-axis to confirm all is working properly.

Details

3:00 Authorize irtc

3:00 Open chamber

3:05 iif_preset, WT10_294, MODE=TRACK, GS=1, ROTMODE=IDLE

3:06 iif_offset, 60, 0 Blank piece of sky for flats. JC an KB take flats in H and Ks. Seems to be too dark, too few counts.

3:12 Preset was canceled because off source for too long.

3:14 iif_preset, WT10_294, MODE=TRACK, GS=1, ROTMODE=IDLE

3:15 iif_offset, 40, 0 Flats.

3:20 Went to zenith so Simone et al could turn off URTH fans

3:31 iif_preset, 'Astar_14', MODE=ACQUIRE, GS=0, ROTMODE=IDLE Rmag=8.77, Hmag=8.0 Steve tweak pointing. Quite a bit of coma. Why is there so much coma (moved secondary -2mm) with the new collimation model.

3:40 Alfio restarted WFS Arbitrator, so we resent the preset. Reference star is near the lower edge of the PISCES image, rather than near the left edge as it was last night. Why? Alfio positioned the W-stage to see the Ref star iif_preset, 'Astar_14', MODE=ACE, GS=3, RS=0, PA=45

3:46 irc_runao

3:50 iif_offset, 10, 0, COORD='DETXY' to put star near center of chip.

3:52 FeII images with Pisces (0014.fits). Bit out of focus, but round. Run focus curve. With improved focus they see that this is a double star (0036.fits). Collect data.

4:05 Steve has been running the DIMM. 1.2" seeing. Guider reports 1.1"

4:38 iif_preset, Astar_15, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • preset failed because the secondary RIPed just before the preset - we wait

4:41 iif_preset, Astar_15, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

4:44 iif_preset, Astar_15, MODE=ACE, GS=1, RS=0, PA=45

4:46 irc_runao SUCCESS!

5:00 iif_preset, Astar_36, MODE=ACQUIRE, GS=0, ROTMODE=IDLE Rmag=8.4 Hmag=7.5

5:04 Steve tweak pointing by 13"

5:05 iif_preset, Astar_36, MODE=ACE, GS=5, RS=0, PA=-90

5:08 irc_runao Success

5:09 iif_offset, -2, 0, COORD='DETXY' Failed on resume AO. Pupils on Pyramid looked OK, but AdSec ripped

5:10 DIMM 1.1"

5:11 iif_preset, Astar_36, MODE=ACE, GS=5, RS=0, PA=-90

5:16 irc_runao - error
 
   ===> Error in irc_runao 
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = AcquireRef[269] complete -  dx:-0.3 dy:-0.031 slNull: f1:Filter #0 f2:Filter #0 freq:990 mag:8.841 nBins:1 nModes:400 r0:0 snMode:0 strehl:0 ttMod:3
        err = Resmsg[2]  = Reading shared buffer masterdiagnostic.R:OPTLOOPBUF (-5001) TIMEOUT_ERROR
        err = Resmsg[3]  = StartAO[270] failed - Reading shared buffer masterdiagnostic.R:OPTLOOPBUF (-5001) TIMEOUT_ERROR
  • the AO loop actually closed BUT iif received the error and turned ON guiding and active optics. This caused the AO loop to fail and rip the AoSec
  • issueTrak 3435

5:24 iif_preset, Astar_36, MODE=ACE, GS=5, RS=0, PA=-90

5:25 irc_runao Success - run focus curves in various bands - DIMM 0.8-1.0''

6:17 iif_preset, Wref28, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - we go to M92 to try AO in a crowded field

6:22 iif_preset, Pisces_9, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

6:23 iif_preset, WT10_330, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - we move to a pointing star newrby because the previous field is too crowded

6:26 iif_preset, Pisces_9, MODE=ACE, GS=19, RS=0, PA=-180

6:30 irc_runao - failed becasue field too crowded and acquisition took too long

6:32 irc_runao

8:00 iif_preset, Pisces_10, MODE=ACE, GS=9, RS=0, PA=90 - offsetting to a star further out at the cluster
  • irc_runao - success

08:28 iif_preset, BS9177, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - seeing limited mode standard stars

08:35 iif_preset, BS9177, MODE=ACTIVE, GS=13, RS=O, PA=-240
  • collimated - taking images in J,Ks,H,H2

9:16 going back to AO mode - iif_preset, Pisces_12, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

9:24 iif_preset, Pisces_12, MODE=ACE, GS=2, RS=0, PA=90
  • irc_runao -failed because the catalog magnitude was 11.3 and the 'real' mag was 14 in the R mag
  • we correct the mag in the catalog and repeat preset

9:28 iif_preset, Pisces_12, MODE=ACE, GS=2, RS=0, PA=90
  • irc_runao

10:44 iif_preset, AO751, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - failed because AGW disconnected...

10:46 iif_preset, AO751, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

10:50 iif_preset, AO751, MODE=ACE, GS=19, RS=0, PA=180
  • could not find the guide star - moved to a nearby pointing star

10:51 iif_preset, WT10_373, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • correct pointing

10:55 iif_preset, AO751, MODE=ACE, GS=19, RS=0, PA=180
  • preset failed agian - check pointing again

10:57 iif_preset, M5_1568, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

10:58 iif_preset, AO751, MODE=ACE, GS=19, RS=0, PA=180
  • irc_runao - failed becuase they had modified the routine - we need to start again from the preset

11:08 iif_preset, AO751, MODE=ACE, GS=19, RS=0, PA=180
  • irc_runao - we cannot get the star because the background is too high

11:14 iif_preset, OC_NGC7243, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • this is a test to see if we can run with an off-axis AO reference star

11:17 iif_preset, OC_NGC7243, MODE=ACE, GS=13, RS=1 , PA=180
  • irc_runao (pisces.20110611.0637.fits) - DIMM 1.1''

11:40 - moving to a star in the west and then we move to a nearby blanck field for flats in J,H,Ks,FeII,H2, Brg

12:30 - End of night.

Sky flat summary

Image   Filter
Pisces.201106.0677-687  J
Pisces.201106.0688-.0698  H
Pisces.201106.0699-0710  Ks
Pisces.201106.0712-0.718 FeII  

-- DougMiller - 11 Jun 2011
Topic revision: r9 - 10 Feb 2012, MichaelWagner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback